A capacitor stores 60micro C charge when connected across a battery .when gap between the plates is filled with dielectric ,charge 120microC flows through the battery . dielectric constant is?

Grade:12

2 Answers

Arun
6 years ago
Dear Ankita
 
We know that the capacitance is directly proportional to the amount of change in the capacitor at a constant voltage. Now since the charge doubled keeping the voltage constant that means the capacitance of the capacitor is doubled after inserting the dielectric material. Now again we know that the capacitance is directly proportional the dielectric constant. Nos since the dielectric constant of the air is almost 1 so the dielectric constant of the material is 2.
 
Regards
Arun (askIITians forum expert)
Ankita Shelke
6 years ago
As initially the capacitor had q=60microCWhen a dielectric was intro of construction say K the amt of charge flowing through emf is 120 So amt of charge on capacitor with dielectric isq`=q+q on emf after dielectric q`=120+60 =180.......(1)Now we know charge on capacitor increases after we put dielectric by K So ,q`=K×qFrom 1 and given K=180/60 =3 So answer is 3